Cole is attempting to lift 150 N. The moment arm of this weight about his elbow joint is 25 cm. The force generated by the elbow flexor muscles is 2000 N. The moment arm of the elbow flexor muscles is 2 cm. Is Cole able to lift the weight with this amount of force in his flexor muscles?
Answer #1

Moment of weight = 150 x 25 = 3750 N.cm


Moment of muscle force (presumeably the maximum she can attain!) = 2000 x 2 = 4000. So she should be able to do the lift with a small margin in reserve.
Answer #2

due to elow flexor = 2000*0.02=40Nm

due to weight = 150*0.25=37.5 Nm

he will be able to lift as applied moment of force is more
Answer #3

Moment of weight = 150 x 25 = 3750 N.cm

Moment of muscle force (presumeably the maximum she can attain!) = 2000 x 2 = 4000. So she should be able to do the lift with a small margin in reserve.